40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 30 15:22:17 2011, Katrin, Update, Green Locking, What happened on green YARM? 
    Reply  Fri Sep 30 16:17:10 2011, kiwamu, Update, Green Locking, things to be done for the Yarm green lock 
Message ID: 5585     Entry time: Fri Sep 30 15:22:17 2011     Reply to this: 5586
Author: Katrin 
Type: Update 
Category: Green Locking 
Subject: What happened on green YARM? 

 

This is a kind of summary of what I have worked on this week.

After all the changes made last week, I could not manage to lock the green light to the cavity, but the PDH error signal looks nicer.....at least something.

 

Alignment of the light to the cavity:

  • DC level of green PD when light is non-resonating 100%
  • DC level of green PD when light resonates 75%
  • --> Not sure if this alignment is good enough
  • In comparision to last week the cavity mirrors seem to move more or my alignment is way worse than last week. The bright spot on ETMY could not be observed for more than let's say a second in the unlocked configuration.

 

Low-pass filter (LPF)

  • The PDH error signal was covered with oscillations of 3.3 kHz, 7.1 kHz and 35 kHz.
  • Measured cut-off frequency of the LPF used so far is 35 kHz
  • Designed and build a new LPF: second order, cut of frequency of 1.1 kHz (this is just the design value, haven't measured that so far)
  • With the new LPF the PDH error signal is free of the above mentioned oscillations.
  • Impedance should be checked

 

PDH error signal

  • Signal-to-noise ratio (SNR) could be improved to values between 7.8 and 11.1 (old SNR was 5 to 7)
  • Looks more like a PDH signal than with the old LPF (now just derivative of the carrier and the first order sidebands show up)
  • Amplitude of the first order sidebands are smaller than the zero order, but are still too high (about 80% of the first order), need to work on the proper value of the LO amplitude an the voltage averager

 

Phase shift between green PD signal and LO

  • Phase shift is about 1MHz
  • Tried to find a capacity that compensates the phase shift. This was not successful since the PD frequency changed every now and than by +/- 20 kHz
ELOG V3.1.3-